### v1

@@ 3.2 Evaluation
Translation quality in this study was assessed using two widely adopted automatic evaluation metrics: SacreBLEU and CHrF++. SacreBLEU Post (2018) provides a standardized and reproducible implementation of the original BLEU metric Papineni et al (2002a), which quantifies n-gram overlap between system-generated translations and human references. Although BLEU-based metrics are prevalent in machine translation evaluation, their emphasis on surface-level lexical matches can limit their effectiveness for morphologically rich languages such as Arabic.
In contrast, CHrF++ Popović (2017) computes F-scores over both character and word-level n-grams, enhancing its sensitivity to morphological variations and orthographic inconsistencies; features that are particularly pronounced in Arabic. As such, CHrF++ is especially well-suited for DA–MSA translation tasks, where minor lexical or syntactic divergences often do not impair semantic equivalence.
Although embedding-based alternatives such as COMET Rei et al (2020) and BERTScore Zhang et al (2019) offer the potential for improved semantic alignment, preliminary experiments that we conducted indicated suboptimal performance for dialectal Arabic. This limitation is likely due to the inadequate representational quality of Arabic embeddings in these models. Consequently, SacreBLEU and CHrF++ were selected for their computational efficiency, interpretability Papineni et al (2002b); Popović (2017), and common use in previous low-resource machine translation researchAbdul-Mageed et al (2024); Habash et al (2024); Mager et al (2023).
These two metrics were employed to provide complementary perspectives: SacreBLEU serves as a precision-oriented measure of lexical fidelity, whereas CHrF++ offers a more balanced assessment of precision and recall at the character level. Both metrics were computed for all model outputs. However, in the subsequent analysis, greater emphasis was placed on CHrF++ scores due to their superior alignment with Arabic’s inflectional and derivational morphology. This evaluation protocol was applied uniformly throughout all experiments to ensure methodological consistency and facilitate comparability.
@@ 3.3.1 Prompting Strategies
The zero shot prompt leverages the pre-trained knowledge of the model to perform translation based solely on an explicit instruction (e.g. “Translate this sentence from the Levantine Dialect to MSA”). Although this approach is efficient and often effective for high-resource language pairs Bouamor et al (2018); Habash (2010), its performance in low-resource scenarios such as DA–MSA translation remains limited. Challenges such as informal orthography, code-switching, and semantic ambiguity frequently reduce translation accuracy. These limitations are consistent with previous findings that highlight the sensitivity of dialect-specific performance to the coverage and composition of the pre-training corpus Sajjad et al (2023); Zhao et al (2023).
@@
Chain-of-Thought (CoT) prompting extends zero-shot prompting by instructing the model to generate intermediate reasoning steps prior to producing the final translation Wei et al (2022). Originally developed for arithmetic and logical inference tasks, CoT was adapted in this study to support dialectal translation. The rationale is that by articulating intermediate reasoning, the model can better resolve idiomatic expressions, lexical ambiguities, and context-dependent phrasing, thus improving translation accuracy.
@@
Few-shot prompting improves the translation quality by including a limited number of DA–MSA example pairs within the prompt. As demonstrated in previous low-resource research Sajjad et al (2023), this strategy enhances generalization by providing concrete examples of the translation task. For each dialect, we curated three representative examples corresponding to the following: conversational, descriptive, and idiomatic sentence types. These categories capture key linguistic features of dialectal Arabic and were selected to expose the model to register variation, structural diversity, and culturally embedded expressions.
@@
To further enhance translation quality, this study introduces Ara-TEaR, an adaptation of the TEaR framework originally proposed by Feng et al. Feng et al (2024). Ara-TEaR incorporates a structured three-stage prompting process: (1) initial translation generation, (2) self-evaluation of the translation against predefined quality criteria, and (3) output refinement based on self-assessment feedback.
Although the original TEaR framework employs Multidimensional Quality Metrics (MQM) Lommel et al (2012, 2024), Ara-TEaR is specifically tailored to the DA–MSA context. It targets two frequent error types: Meaning Transfer errors, in which the intended meaning is not adequately preserved, and Adaptation errors, which occur when the style fails to shift appropriately from dialectal Arabic to formal MSA. These error types were identified through manual analysis of the outputs generated by three baseline models: GPT-3.5, Jais, and NLLB-200. A visual overview of the Ara-TEaR workflow is presented in Section  9.
@@ 3.3.2 Model and Dataset Selection
In our model prompting experiments, we evaluate seven LLMs to examine how prompting effectiveness varies across architectures and parameter scales. Previous research has shown that model performance is influenced by both architectural design and the composition of pre-training data Zhao et al (2023). As shown in Table 3, the selected models span a range of sizes and architectures, allowing a comprehensive analysis of their respective strengths and limitations in the DA–MSA translation task.
@@
GPT-4o Approximately 1.8T Exploding Topics (2025)
@@
Table 3: Overview of evaluated models and their parameter sizes, highlighting architectural diversity in the DA–MSA translation experiments.
Our model selection was intended to reflect a diverse spectrum of capabilities and design philosophies. GPT-3.5 was included as a widely adopted general-purpose baseline, while GPT-4o and DeepSeek-V3 represent recent advancements in multilingual LLMs. LLaMA 3.3-70B-Instruct and LLaMA 3.1-405B facilitate analysis of instruction tuning and scaling in open-source systems. Gemma-27B was chosen for its balance between computational efficiency and observed performance.
As no parameter updates were performed during prompting, training data was not required. Instead, evaluation relied solely on the T1 test set, derived from the MADAR corpus, comprising 600 examples. While this test set is smaller than our T2 test set (1,200 examples), which is used in the model fine-tuning experiment, T1’s size was also chosen with practical considerations in mind. Given the prohibitive token-based usage costs associated with API access to these models, T1 was designed to be small enough to allow us to complete all our experiments, which involve extensive evaluation across different models and prompting strategies. This setup supported our systematic and resource-conscious evaluation across model variants.
@@ 3.3.3 Implementation Details
The prompting experiments were conducted via API access to each of the selected language models. Specifically, the OpenAI API was utilized for GPT-3.5 and GPT-4o, the Anthropic API for Claude 3.5 Sonnet, and the LLAMA API for the open-source models: LLaMA 3.3-70B-Instruct, LLaMA 3.1-405B, Gemma-27B, and DeepSeek-V3.
@@ 3.4.1 Model and Dataset Selection
The Gemma-2-9B model was selected as the primary architecture for the majority of fine-tuning experiments due to its balance between computational efficiency and translation quality. Its strong performance in initial zero-shot prompting trials, in addition to its effectiveness in multilingual tasks, particularly for non-English language pairs, make it suitable for DA–MSA translation. With 9 billion parameters, Gemma-2-9B provides sufficient capacity to capture dialectal variation while remaining computationally feasible for fine-tuning. Additional features of this model, such as native support for 4-bit quantization and Low-Rank Adaptation (LoRA), further reduced memory requirements, enabling efficient training under low-resource conditions.
Fine-tuning experiments utilized the three dataset compositions: D1, D2, and D3. All data were formatted using an Alpaca-style instruction template Taori et al (2023), adapted for DA–MSA translation:
@@
This format capitalizes on the instruction-following capabilities of modern LLMs while omitting explicit dialect labels, thus encouraging implicit dialect identification. To mitigate any training data biases, all datasets were randomly shuffled.
Evaluation primarily relied on the T2 test set, which comprises 1,200 examples (400 per dialect) drawn from both MADAR (70%) and Dial2MSA-Verified (30%). The heavier weighting toward MADAR reflects its greater reliability due to professional translation, while the inclusion of Dial2MSA-Verified data introduces additional lexical and stylistic diversity from social media contexts. During initial hyperparameter optimization, the smaller T1 set (600 examples) was employed as a development set to reduce evaluation costs. The size of our dataset aligns with common DA-MSA shared tasks, such as NADI-2024 Abdul-Mageed et al (2024), where the test dataset size was 400 examples per dialect, and OSACT-2024 Habash et al (2024), where the dataset sizes for the Egyptian, Levantine, and Gulf dialects were 314, 568, and 586 examples, respectively.
@@ 3.4.2 Fine-Tuning Configuration
The experiments were conducted using A100 GPUs on Google Colab Pro. Low-Rank Adaptation (LoRA) was applied with a rank of 16 and an alpha parameter of 16. The key training settings included a batch size of 16 and gradient accumulation steps of 2. The AdamW optimizer was employed in 8-bit mode, with a weight decay coefficient of 0.01. A linear learning rate scheduler was utilized, and mixed precision training (FP16/BF16) was adopted to improve computational efficiency.
@@ 3.4.3 Experimental Framework
A series of ablation experiments was conducted to systematically assess factors influencing the performance of DA–MSA translation models, including model architecture, data strategy, and training hyperparameters. Beginning with a baseline configuration (Experiment 0), each subsequent experiment introduced a controlled modification to isolate the effect of a single variable. This structured approach facilitated reproducible comparisons and yielded targeted insights for optimizing DA–MSA translation in low-resource settings.
Experiment 0: Baseline Performance
The primary objective of this experiment was to establish a baseline for DA–MSA translation performance using untuned models. This baseline served as a reference point for evaluating the impact of subsequent fine-tuning experiments. The evaluation focused on the Gemma-2-9B model in two configurations: the base model (without instruction tuning) and an instruction-tuned variant. Both configurations were evaluated under a zero-shot inference setting, without any fine-tuning. All evaluations were conducted on the designated test set, following the setup described earlier.
Experiment 1: Hyperparameter Optimization
This experiment aimed to identify optimal values for the learning rate and the number of training epochs to balance model performance and training stability. These hyperparameters were prioritized due to their substantial impact on convergence behavior and overfitting. The D1 dataset (MADAR: 43,000 samples) was selected for its sufficient size to support meaningful training while avoiding the complexity and noise associated with larger datasets. The Gemma-2-9B model was employed with consistent architectural and training configurations, while systematically varying the learning rate (1e-4 vs. 5e-5) and the number of training epochs (1 vs. 3).
Experiment 2: Joint vs. Dialect-Specific Models
@@
    Joint Model: Trained on the complete D1 dataset, covering all three dialects (LEV, GLF, EGY).
@@
    EGY-Specific Model: Trained exclusively on the Egyptian subset of D1.
Experiment 3: Data Scaling Effects
This experiment evaluated the impact of increasing training data volume and introducing domain diversity on translation performance. The Gemma-2-9B model was fine-tuned using the D2 dataset (MADAR + Dial2MSA), expanding the training set from 43,000 to 62,775 examples. All training configurations were kept consistent with those employed in the previous experiment.
Experiment 4: Data Augmentation
This experiment aimed to evaluate whether incorporating synthetic data generated through MSA paraphrasing could improve DA–MSA translation quality. The training dataset D3, combined D2 (comprising 62,775 samples) with two additional components: 43,000 paraphrased MSA sentences generated using GPT-4o, and 5,000 same-source-target MSA–MSA pairs extracted from the UFAL corpus. This resulted in a total of 110,775 training samples. Two model variants, Gemma-9B and Gemma-27B, were fine-tuned on this dataset to investigate whether model size influences the effectiveness of the data augmentation strategy. GPT-4o was chosen for paraphrase generation due to its strong zero-shot performance and relatively low cost.
Experiment 5: Model Architecture Comparison
This experiment evaluated DA–MSA translation across various model architectures and sizes. Its objectives are to determine whether smaller fine-tuned models could match the performance of larger zero-shot models, identify architecture-specific strengths, and provide practical insights for deployment in resource-constrained settings. Models were selected based on their architectural diversity, parameter size, and multilingual capabilities. Dataset D2 (MADAR + Dial2MSA), identified as the best-performing configuration in Experiment 3, was used for training. All models were fine-tuned using identical hyperparameters (learning rate = 5e-5 and one epoch). Table 4 presents the fine-tuned and zero-shot prompted models evaluated in this experiment.
@@
Table 4: Fine-tuned and zero-shot models used in Experiment 5.
Experiment 6: Quantization Impact Analysis
This experiment assessed the trade-offs between translation quality and computational efficiency resulting from applying 4-bit quantization. The objective is to validate the use of quantization in our prior experiments and examine its viability for real-world deployment. The LLaMA-3.2-3B-Instruct model was fine-tuned on Dataset D2 using two precision settings: full-precision (16-bit), which served as the baseline, and 4-bit quantization, which was used to explore potential resource savings and quality retention. This model was selected due to its smaller size, which enabled fine-tuning under both configurations given the available hardware resources.
Experiment 7: Active Learning Approach
This experiment investigated the effect of selective fine-tuning using linguistically challenging examples to determine whether emphasizing complex sentences could improve generalization in dialect-to-MSA translation. The Gemma-9B model, previously fine-tuned on Dataset D2, was further fine-tuned using a subset consisting of the longest 30% of sentences from D2. This subset was selected based on the assumption that longer sentences are more likely to exhibit greater syntactic and semantic complexity.
4 Results and Discussion
4.1 Training-free prompting approach for DA–MSA translation

### v2

@@ 3.2 Evaluation
Translation quality in this study was assessed using two widely adopted automatic evaluation metrics: SacreBLEU and chrF++. SacreBLEU (Post, 2018) provides a standardized and reproducible implementation of the original BLEU metric (Papineni et al, 2002), which quantifies n-gram overlap between system-generated translations and references. Although BLEU-based metrics are prevalent in machine translation evaluation, their emphasis on surface-level lexical matches can limit their effectiveness for morphologically rich languages such as Arabic (Han et al, 2021).
In contrast, chrF++ (Popović, 2017) computes F-scores over both character and word-level n-grams, enhancing its sensitivity to morphological variations and orthographic inconsistencies; features that are particularly pronounced in Arabic. As such, chrF++ is especially well-suited for DA–MSA translation tasks, where minor lexical or syntactic divergences often do not impair semantic equivalence.
Although embedding-based alternatives such as COMET (Rei et al, 2020) and BERTScore (Zhang et al, 2019) offer the potential for improved semantic alignment, preliminary experiments that we conducted indicated suboptimal performance for dialectal Arabic. This limitation is likely due to the inadequate representational quality of Arabic embeddings in these models. While the LLM-as-a-judge paradigm has recently gained popularity in machine translation (MT) evaluation, where an LLM is prompted to assign scores reflecting the adequacy or correctness of a translation (Bavaresco et al, 2025), its use in this study was infeasible due to the prohibitive cost of large-scale prompting. Furthermore, such approaches offer limited interpretability, as the decision-making process behind the assigned scores remains opaque. Consequently, SacreBLEU and chrF++ were selected for their computational efficiency, interpretability (Papineni et al, 2002; Popović, 2017), and common use in previous low-resource machine translation research(Abdul-Mageed et al, 2024; Al-Khalifa et al, 2024; Mager et al, 2023).
These two metrics were employed to provide complementary perspectives: SacreBLEU serves as a precision-oriented measure of lexical fidelity, whereas chrF++ offers a more balanced assessment of precision and recall at the character level. Both metrics were computed for all model outputs. However, in all subsequent results analysis, greater emphasis was placed on chrF++ scores due to their superior suitability for Arabic’s inflectional and derivational morphology. This evaluation protocol was applied uniformly throughout all experiments to ensure methodological consistency and facilitate comparability.
@@ 3.3.1 Prompting Strategies
The Zero-Shot prompt leverages the pre-trained knowledge of the model to perform translation based solely on an explicit instruction (e.g. “Translate this sentence from the Levantine Dialect to MSA”). Although this approach is efficient and often effective for high-resource language pairs (Johnson et al, 2017; Brown et al, 2020), its performance in low-resource scenarios such as DA–MSA translation remains limited. Challenges such as informal orthography, code-switching, and semantic ambiguity frequently reduce translation accuracy. These limitations are consistent with previous findings that highlight the sensitivity of dialect-specific performance to the coverage and composition of the pre-training corpus (Sajjad et al, 2023; Zhao et al, 2023).
@@
Chain-of-Thought (CoT) prompting extends zero-shot prompting by instructing the model to generate intermediate reasoning steps prior to producing the final translation (Wei et al, 2022). Originally developed for arithmetic and logical inference tasks, CoT was adapted in this study to support dialectal translation. The rationale is that by articulating intermediate reasoning, the model can better resolve idiomatic expressions, lexical ambiguities, and context-dependent phrasing, thus improving translation accuracy.
@@
Few-shot prompting improves the translation quality by including a limited number of DA–MSA example pairs within the prompt. As demonstrated in previous low-resource research (Sajjad et al, 2023), this strategy enhances generalization by providing concrete examples of the translation task. For each dialect, we curated three representative examples corresponding to the following: conversational, descriptive, and idiomatic sentence types. These categories capture key linguistic features of dialectal Arabic and were selected to expose the model to register variation, structural diversity, and culturally embedded expressions.
@@
To further enhance translation quality, we introduce Ara-TEaR, an adaptation of the TEaR framework originally proposed by  Feng et al (2025). Ara-TEaR incorporates a structured three-stage prompting process: (1) initial translation generation, (2) self-evaluation of the translation against predefined quality criteria, and (3) output refinement based on self-assessment feedback.
While the original TEaR framework employs Multidimensional Quality Metrics (MQM) (Lommel et al, 2012, 2024), Ara-TEaR is specifically tailored to the DA–MSA context. It targets two frequent error types: Meaning Transfer errors, in which the intended meaning is not adequately preserved, and Adaptation errors, which occur when the style fails to shift appropriately from dialectal Arabic to formal MSA. These error types were identified through manual analysis of the outputs generated by three baseline models: GPT-3.5, Jais, and NLLB-200. A visual overview of the Ara-TEaR workflow is presented in Figure 7.
@@ 3.3.2 Model and Dataset Selection
In our model prompting experiments, we evaluate six LLMs to examine how prompting effectiveness varies across architectures and parameter scales. Previous research has shown that model performance is influenced by both architectural design (Vaswani et al, 2017) and the composition of pre-training data (Zhao et al, 2023). As shown in Table 3, the selected models span a range of sizes and architectures, allowing a comprehensive analysis of their respective strengths and limitations in the DA–MSA translation task.
@@
GPT-4o Approximately 1.8T (Exploding Topics, 2025)
@@
Table 3: Overview of evaluated models and their parameter sizes, highlighting architectural diversity in the DA–MSA translation experiments
Our model selection was intended to reflect a diverse spectrum of capabilities and design philosophies. GPT-3.5 was included as a widely adopted general-purpose baseline, while GPT-4o and DeepSeek-V3 represent recent advancements in multilingual LLMs. LaMA 3.3-70B-Instruct and LLaMA 3.1-405B facilitate analysis of instruction tuning and scaling in open-source systems. Gemma-27B was chosen for its computational efficiency and observed good performance during manual testing.
As no parameter updates were performed during prompting, training data was not required. Instead, evaluation relied solely on the Small-Test test set, derived from the MADAR corpus, comprising 600 examples. While this test set is smaller than our Large-Test dataset (1,200 examples), which is used in the model fine-tuning experiment, Small-Test dataset’s size was also chosen with practical considerations in mind. Given the prohibitive token-based usage costs associated with API access to these models, Small-Test was designed to be small enough to allow us to complete all our experiments, which involve extensive evaluation across different models and prompting strategies. This setup supported our systematic and resource-conscious evaluation across the different model variants.
@@ 3.3.3 Implementation Details
The prompting experiments were conducted via API access to each of the selected models. Specifically, the OpenAI API was utilized for GPT-3.5 and GPT-4o, the Anthropic API for Claude 3.5 Sonnet, and the LLAMA API for the open-source models: LLaMA 3.3-70B-Instruct, LLaMA 3.1-405B, Gemma-27B, and DeepSeek-V3.
@@ 3.4.1 Model and Dataset Selection
The Gemma-2-9B model was selected as the primary architecture for the majority of fine-tuning experiments due to its balance between computational efficiency and translation quality. In addition to its effectiveness in multilingual tasks, particularly for non-English language pairs, Gemma-2-9B demonstrates strong translation capabilities across many languages (Cui et al, 2025), making it suitable for DA–MSA translation. With 9 billion parameters, it provides sufficient capacity to capture dialectal variation while remaining computationally feasible for fine-tuning. Additional features of this model, such as native support for 4-bit quantization and LoRA, further reduced memory requirements, enabling efficient training under low-resource conditions.
Fine-tuning experiments utilized the three dataset compositions: Minimal-Gold , Extended-Gold, and Extended-Gold-Silver. All data were formatted using an Alpaca-style instruction template (Taori et al, 2023), adapted for DA–MSA translation:
@@
This format capitalizes on the instruction-following capabilities of modern LLMs while omitting explicit dialect labels, thus encouraging implicit dialect identification. To mitigate any training data biases, all training datasets were randomly shuffled.
Evaluation primarily relied on the Large-Test test set, which comprises 1,200 examples (400 per dialect) drawn from both MADAR (70%) and Dial2MSA-Verified (30%). The heavier weighting toward MADAR reflects its greater reliability due to professional translation, while the inclusion of Dial2MSA-Verified data introduces additional lexical and stylistic diversity from social media contexts. During initial hyperparameter optimization, the smaller Small-Test set (600 examples) was employed as a development set to reduce evaluation costs. The size of our dataset aligns with common DA-MSA shared tasks, such as NADI-2024 (Abdul-Mageed et al, 2024), where the test dataset size was 400 examples per dialect, and OSACT-2024 (Al-Khalifa et al, 2024), where the dataset sizes for the Egyptian, Levantine, and Gulf dialects were 314, 568, and 586 examples, respectively.
@@ 3.4.2 Fine-Tuning Configuration
The experiments were conducted using A100 GPUs on Google Colab Pro. LoRA was applied with a rank and an alpha parameter of 16. The key training settings included a batch size of 16 and gradient accumulation steps of 2. The AdamW optimizer was employed in 8-bit mode, with a weight decay coefficient of 0.01. A linear learning rate scheduler was utilized, and mixed precision training (FP16/BF16) was adopted to improve computational efficiency.
@@ 3.4.3 Experimental Framework
A series of ablation experiments (8 in total) was conducted to systematically assess factors influencing the performance of our models, including model architecture, data strategy, and training hyperparameters. Beginning with a baseline configuration, each subsequent experiment introduced a controlled modification to isolate the effect of a single variable. This structured approach facilitated reproducible comparisons and yielded targeted insights for optimizing DA–MSA translation in low-resource settings. The 8 ablation experiments are named as below and will be explained in details.
  • Base: Baseline Performance (v0)
  • HypperParam: Hyperparameter Optimization (v1)
  • Join-vs-Spec: Joint vs. Dialect-Specific Models (v2)
  • DataScale: Data Scaling Effect (v3)
  • Augmentation: Data Augmentation Effect (v4)
  • Architecture: Model Architecture Comparison (v5)
  • Quantiz: Quantization Impact (v6)
  • ActiveLearn: Active Learning Approach (v7)
Base: Baseline Performance (v0)
The primary objective of this experiment was to establish a baseline for DA–MSA translation performance using an untuned model. This baseline served as a reference point for evaluating the impact of subsequent fine-tuning experiments. The evaluation focused on the Gemma-2-9B model in two configurations: the base model (without instruction tuning) and an instruction-tuned variant. Both configurations were evaluated under a zero-shot inference setting, without any fine-tuning. All evaluations were conducted on the designated test set, following the setup described earlier.
HypperParam: Hyperparameter Optimization (v1)
This experiment aimed to identify optimal values for the learning rate and the number of training epochs to balance model performance and training stability. These hyperparameters were prioritized due to their substantial impact on convergence behavior and overfitting. The Minimal-Gold dataset (MADAR: 43,000 samples) was selected for its sufficient size to support meaningful training while avoiding the complexity and noise associated with larger datasets. The Gemma-2-9B model was employed with consistent architectural and training configurations, while systematically varying the learning rate (1e-4 vs. 5e-5) and the number of training epochs (1 vs. 3).
Join-vs-Spec: Joint vs. Dialect-Specific Models (v2)
@@
    Joint Model: Trained on the complete Minimal-Gold dataset, covering all three dialects (LEV, GLF, EGY).
@@
    EGY-Specific Model: Trained exclusively on the Egyptian subset of the Minimal-Gold dataset.
DataScale: Data Scaling Effect (v3)
This experiment evaluated the impact of increasing training data volume and introducing domain diversity on translation performance. The Gemma-2-9B model was fine-tuned using the Extended-Gold dataset (MADAR + Dial2MSA-Verified), expanding the training set from 43,000 to 62,775 examples. All training configurations were kept consistent with those employed in the previous experiment.
Augmentation: Data Augmentation Effect (v4)
This experiment aimed to evaluate whether incorporating synthetic data generated through MSA paraphrasing could improve DA–MSA translation quality. The Extended-Gold-Silver training dataset, combined the Extended-Gold dataset (comprising 62,775 samples) with two additional components: 43,000 paraphrased MSA sentences generated using GPT-4o, and 5,000 same-source-target MSA–MSA pairs extracted from the UFAL corpus. This resulted in a total of 110,775 training samples. Two model variants, Gemma-9B and Gemma-27B, were fine-tuned on this dataset to investigate whether model size influences the effectiveness of data augmentation. GPT-4o was chosen for paraphrase generation due to its strong zero-shot performance and relatively low cost.
Architecture: Model Architecture Comparison (v5)
This experiment evaluated DA–MSA translation across various model architectures and sizes. Its objectives are to determine whether smaller fine-tuned models could match the performance of larger zero-shot models, identify architecture-specific strengths, and provide practical insights for deployment in resource-constrained settings. Models were selected based on their architectural diversity, parameter size, and multilingual capabilities. The Extended-Gold dataset (MADAR + Dial2MSA-Verified ), identified as the best-performing configuration in the DataScale experiment, was used for training. All models were fine-tuned using identical hyperparameters (learning rate = 5e-5 and one epoch). Table 4 presents the fine-tuned and zero-shot prompted models evaluated in this experiment.
@@
Table 4: Fine-tuned and zero-shot models evaluated in the Architecture experiment
Quantiz: Quantization Impact (v6)
This experiment assessed the trade-offs between translation quality and computational efficiency resulting from applying 4-bit quantization. The objective is to validate the use of quantization in our prior experiments and examine its viability for real-world deployment. The LLaMA-3.2-3B-Instruct model was fine-tuned on the Extended-Gold dataset using two precision settings: full-precision (16-bit), which served as the baseline, and 4-bit quantization, which was used to explore potential resource savings and quality retention. This model was selected due to its smaller size, which enabled fine-tuning under both configurations given the available hardware resources.
ActiveLearn: Active Learning Approach (v7)
This experiment investigated the effect of selective fine-tuning using linguistically challenging examples to determine whether emphasizing complex sentences could improve generalization in dialect-to-MSA translation. The Gemma-9B model, previously fine-tuned on the Extended-Gold dataset, was further fine-tuned using a subset consisting of the longest 30% of sentences from the Extended-Gold dataset. This subset was selected based on the assumption that longer sentences are more likely to exhibit greater syntactic and semantic complexity.
4 Experimental Results
4.1 Evaluation of the Training-free prompting approach for DA–MSA translation